17 You may want to reference (Page) section 16.2 while completing this problem. Use the Henderson-Hasselbalch...

17

You may want to reference (Page) section 16.2 while completing this problem.

Use the Henderson-Hasselbalch equation to calculate the pH of each solution:

Part A

a solution that is 0.170 M in HC2H3O2 and 0.125 M in KC2H3O2

Express your answer using two decimal places.

pH=

Part B

a solution that is 0.205 M in CH3NH2 and 0.135 M in CH3NH3Br

Express your answer using two decimal places.

pH=
Answer #1

Part A) Given, a buffer solution that is 0.170 M in HC2H3O2 and 0.125 M in KC2H3O2

We know the Ka value for acetic acid = 1.8 x 10-5

Thus, pKa = -logKa

pKa = -log(1.8 x 10-5)

pKa = 4.7447

Now, the equilibrium representation of this solution is,

HC2H3O2(aq) + H2O(l) C2H3O2-(aq) + H3O+(aq)

We know, the Henderson-Hasselbalch equation,

pH = pKa + log [C.base / Acid]

Substituting the known values,

pH = 4.7447 + log [0.125 / 0.170]

pH = 4.7447 - 0.1335

pH = 4.61

Part B)

Given, a buffer solution that is 0.205 M in CH3NH2 and 0.135 M in CH3NH3Br

We know, the Kb value for CH3NH2 = 4.4 x 10-4

Thus, pKb = -logKb

pKb = -log(4.4 x 10-4)

pKb = 3.36

Now, the equilibrium representation of this solution is,

CH3NH2(aq) + H2O(l) CH3NH3+(aq) + OH-(aq)

We know, the Henderson-Hasselbalch equation,

pOH = pKb + log [C.Acid / Base]

Substituting the known values,

pOH = 3.36 + log [0.135 / 0.205]

pOH = 3.36 - 0.1814

pOH = 3.1751

Now, We know,

pH + pOH = 14

pH + 3.1751 = 14

pH = 10.82